OpenStudy (anonymous):

f(x) = x3 + 4x2 - x - 4 Part A: What are the factors of f(x)? Show your work. (3 points) Part B: What are the zeros of f(x)? Show your work. (2 points) Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (zale101):

Do you recall group factoring?

OpenStudy (anonymous):

No :(

OpenStudy (zale101):

f(x) = x3 + 4x2 - x - 4 f(x) = (x^3 + 4x^2) -( x + 4 ) Can you factor each parenthesis now ?

OpenStudy (anonymous):

Im not sure im honestly so new to this :(

OpenStudy (zale101):

f(x) = (x^3 + 4x^2) -( x + 4 ) x-4 became x+4 because i factored a negative one just so you're not confused

OpenStudy (zale101):

https://www.mathsisfun.com/algebra/factoring.html Hope this could help! Let us do some brain exercise here. Let's say y=x^3+2x y=(x^3+2x) What is the common factor?

OpenStudy (triciaal):